# Interpretable Reward Model via Sparse Autoencoder

### v4

@@ Introduction
We train SARM exclusively on the Skywork-Reward-Preference-80K-v0.2 dataset (SkyworkReward) for 3 epochs. As illustrated in Figure 2, we discard all layers after depth $l$, freeze the parameters of the pretrained SAE encoder, and train only the first $l$ layers of the backbone language model along with the final linear head. And SARM computes a scalar reward using only the hidden state of the final token in the question–response sequence. We set the global batch size to $8$ and the learning rate to $4\times 10^{-6}$.
@@
(a) Manipulating a safety-related feature shifts reward distribution on the safety-related dataset.
(b) The safety-unrelated dataset shows no significant change after safety feature manipulating.
Figure 3: Effect of manipulating a safety-related feature on reward distribution. (a) Manipulating a safety-related feature leads to a clear rightward shift in the reward distribution, indicating successful preference manipulation. (b) The distribution remains largely unchanged on the safety-unrelated dataset, suggesting that the manipulating effect does not influence irrelevant attributes.
Since SARM directly attributes reward scores to interpretable features, it is natural to ask whether manipulating these features enables control over the reward model’s preferences. To investigate this, we focus specifically on safety as an representative use case.
To pinpoint safety-related features, we adopt the approach described in (safer). We measure the activation differences of each feature between chosen and rejected responses using a safety-focused preference dataset, specifically the Safety subset of Reward Bench 2 (RewardBench). The overall activation difference between for each feature is denoted as $s_{i}$. In this case, we select the feature with the highest $s_{i}$ as our target for intervention.
Due to the Top-$k$ activation function, SARM attributes the reward score not across all $M$ features, but only across the $k$ activated features. Our objective is therefore to manipulate preference of SARM using the target feature. Specifically, we multiply the weight corresponding to the selected feature with the largest $s_{i}$ in the linear head by a constant. As shown in Figure 3, manipulating this feature causes a clear rightward shift in the reward scores on the safety-related dataset (cf. Figure 3(a)), indicating that SARM assigns higher rewards to safer responses as intended. In contrast, when applied to a safety-unrelated dataset, the reward distribution remains largely unchanged (cf. Figure 3(b)), suggesting that the intervention selectively affects relevant contexts and does not influence unrelated attributes. This result demonstrates that the learned features are not only interpretable but also causally controllable, allowing precise manipulation of reward model preferences.
@@
SARM-1B 1.5B 62.5 55.6 35.6 60.7 84.9 82.4 56.0
SARM-2B 2.4B 64.2 58.6 34.4 62.8 87.3 86.3 55.6
SARM-4B 4.2B 73.6 68.5 42.5 63.9 91.3 96.0 79.6
@@
Notably, our SARM-4B achieves the best overall score across all models (73.6), outperforming both open- and closed-source baselines, including models with significantly larger parameter counts such as Llama-3.1-Tulu-3-70B-SFT. This result demonstrates that incorporating interpretability via sparse feature modeling does not degrade alignment quality. Instead, SARM benefits from interpretable and controllable reward representation, while maintaining strong generalization across diverse evaluation axes. Furthermore, SARM-1B and SARM-2B, both derived from a 3B backbone, perform competitively despite using significantly fewer parameters than most open-source baselines. This suggests that even at smaller scales, interpretable reward models can retain strong alignment performance with proper architectural design.
@@
SAE Encoder with Random Init (4+0.2) B 68.4 65.7 38.4 64.5 88.9 88.2 64.9
Token-level SAE Pretraining (4+0.2) B 71.5 68.0 40.6 62.3 92.9 92.5 72.5
SARM-4B (4+0.2) B 73.6 68.5 42.5 63.9 91.3 96.0 79.6
@@ Conclusion
Layer 7 (0.75+0.15) B 36.21 37.26 31.25 46.72 43.77 38.78 19.51
Layer 10 (1+0.15) B 53.55 49.89 31.87 49.18 82.88 67.47 38.70
Layer 14 (SARM) (1.5+0.15) B 62.52 55.57 35.62 60.65 84.88 82.42 55.97
Layer 21 (2.25+0.15) B 64.17 58.63 34.37 62.84 87.33 86.26 55.60
Layer 28 (3+0.15) B 62.94 57.68 36.25 61.20 85.50 80.80 56.19

### v5

@@ Introduction
We train SARM exclusively on the Skywork-Reward-Preference-80K-v0.2 dataset (SkyworkReward) for 3 epochs. As illustrated in Figure 2, we discard all layers after depth $l$, freeze the parameters of the pretrained SAE encoder, and train only the first $l$ layers of the backbone language model along with the final linear head. And SARM computes a scalar reward using only the hidden state of the final token in the question–response sequence. We set the global batch size to $512$ and the learning rate to $4\times 10^{-6}$.
@@
(a) Manipulating a safety-related feature shifts reward distribution on the target dataset $T$.
(b) The complement of target dataset $C$ shows no significant change after safety feature manipulating.
Figure 3: Effect of manipulating a safety-related feature on reward distribution. (a) Manipulating a safety-related feature leads to a clear rightward shift in the reward distribution, indicating successful preference manipulation. (b) The distribution remains largely unchanged on the complement of target datasets, suggesting that the manipulating effect does not influence irrelevant attributes.
Since SARM directly attributes reward scores to interpretable features, it is natural to investigate whether manipulating these features enables control over the reward model’s preferences. To explore this, we focus specifically on safety as a representative use case.
To identify safety-related features, we compute a score $s_{i}$ for each feature following the method in (safer), where $s_{i}$ quantifies how frequently the $i$-th feature activates on contexts positively correlated with safety. This correlation is estimated from activation differences between chosen and rejected responses within a safety-focused preference dataset (the safety subset of RM-Bench (RM_Bench)). The feature with the highest $s_{i}$ is selected for intervention. We then multiply the $i$-th value-head parameter by a constant and evaluate its effect on RewardBench 2 (RewardBench_2). Due to the Top-$K$ activation, only active features contribute to the reward, so the intervention affects scores only when the target feature is activated. Ideally, this shifts the reward distribution of safety-related data rightward while leaving unrelated data largely unchanged.
For this evaluation, we reformulate RewardBench 2 (RewardBench_2), whose instances are preference triplets $(x,y^{+},y^{-})$. We convert each triplet into individual query–response pairs $(x,y)$ and partition them into two disjoint sets. Target set $T$: pairs where the query $x$ belongs to the safety subset and the response is the chosen one. Complement set $C$: all remaining pairs, including (i) safety queries paired with the rejected response and (ii) non-safety queries paired with either the chosen or rejected response.
As shown in Figure 3, this intervention causes a clear rightward shift in the reward score distribution for the Target Set $T$ (cf. Figure 3(a)). This indicates that SARM assigns higher rewards to safer responses, as intended. Conversely, the reward distribution for the Complement Set $C$ remains largely unchanged (cf. Figure 3(b)), suggesting the intervention selectively impacts relevant contexts without affecting unrelated attributes. This result demonstrates that the learned features are not only interpretable but also causally controllable, allowing precise manipulation of reward model preferences.
@@
SARM-2B 2.0B 62.5 55.6 35.6 60.7 84.9 82.4 56.0
SARM-3B 2.7B 64.2 58.6 34.4 62.8 87.3 86.3 55.6
SARM-4B 4.3B 73.6 68.5 42.5 63.9 91.3 96.0 79.6
@@
Notably, our SARM-4B achieves the best overall score across all models (73.6), outperforming both open- and closed-source baselines, including models with significantly larger parameter counts such as Llama-3.1-Tulu-3-70B-SFT. This result demonstrates that incorporating interpretability via sparse feature modeling does not degrade alignment quality. Instead, SARM benefits from interpretable and controllable reward representation, while maintaining strong generalization across diverse evaluation axes. Furthermore, SARM-2B and SARM-3B, both derived from a 3B backbone, perform competitively despite using significantly fewer parameters than most open-source baselines. This suggests that even at smaller scales, interpretable reward models can retain strong alignment performance with proper architectural design.
@@
SAE Encoder with Random Init (4+0.3) B 68.4 65.7 38.4 64.5 88.9 88.2 64.9
Token-level SAE Pretraining (4+0.3) B 71.5 68.0 40.6 62.3 92.9 92.5 72.5
SARM-4B (4+0.3) B 73.6 68.5 42.5 63.9 91.3 96.0 79.6
@@ Conclusion
Acknowledgments
This research is supported by the National Science and Technology Major Project (2023ZD0121102). This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Acknowledgments
Layer 7 (1.1+0.15) B 36.21 37.26 31.25 46.72 43.77 38.78 19.51
Layer 10 (1.4+0.15) B 53.55 49.89 31.87 49.18 82.88 67.47 38.70
Layer 14 (SARM) (1.8+0.15) B 62.52 55.57 35.62 60.65 84.88 82.42 55.97
Layer 21 (2.5+0.15) B 64.17 58.63 34.37 62.84 87.33 86.26 55.60
Layer 28 (3.2+0.15) B 62.94 57.68 36.25 61.20 85.50 80.80 56.19